### v2

# MARS-M: When Variance Reduction Meets Matrices

Abstract
Matrix-based preconditioned optimizers, such as Muon, have recently been shown to be more efficient than scalar-based optimizers for training large-scale neural networks, including large language models (LLMs). On the other hand, recent benchmarks on optimizers for LLM pre-training have demonstrated that variance-reduction techniques such as MARS can achieve substantial speedups over standard optimizers that do not employ variance reduction. In this paper, to achieve the best of both worlds, we introduce MARS-M, a new optimizer that integrates the variance reduction technique in MARS with Muon. Under standard regularity conditions, we prove that Muon-M converges to a first-order stationary point at a rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, which improves upon $\widetilde{\mathcal{O}}(T^{-1/4})$ rate attained by Muon. Our empirical results on language modeling and computer vision tasks demonstrate that MARS-M consistently yields lower losses and improved performance across various downstream benchmarks. The implementation of MARS-M is available at https://github.com/AGI-Arena/MARS/tree/main/MARS_M.
1 Introduction
The development of preconditioned gradient methods, such as AdamW (Loshchilov and Hutter, 2019), AdaFactor (Shazeer and Stern, 2018) and Lion (Chen et al., 2023), plays an important role in the advance of large-scale deep learning. A lot of renowned large language models (LLMs), including ChatGPT (OpenAI, 2023), LLaMA-3 (Dubey et al., 2024) and DeepSeek-R1 (Guo et al., 2025) are trained with adaptive gradient methods such as Adam (Kingma and Ba, 2015) and AdamW. Recently, matrix-based preconditioned optimization methods, such as Shampoo (Gupta et al., 2018), SOAP (Vyas et al., 2024) and Muon (Jordan et al., 2024; Liu et al., 2025), have been introduced to speed up the training of large models like Kimi K2 (Team et al., 2025) and GLM-4.5 (Z.ai, 2025). Unlike vector-based methods, matrix-based approaches operate directly on the model parameter matrices without flattening them, thereby preserving their inherent 2D structure and the underlying optimization geometry.
On the other hand, stochastic optimization methods are often hindered by high variance in the stochastic gradients during training, which can slow convergence and degrade stability. To address this issue, numerous variance reduction techniques have been proposed, including SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013), SARAH (Nguyen et al., 2017a, b), SPIDER (Fang et al., 2018), SNVRG (Zhou et al., 2020) and STORM (Cutkosky and Orabona, 2019), to mention a few. However, these methods have seen limited success in training large-scale deep neural networks, largely due to incompatibilities with the practice of deep learning and neural network architectures (Defazio and Bottou, 2019). To make variance reduction work for training large-scale deep neural networks and LLMs, MARS (Yuan et al., 2025) was recently proposed. It introduces a scaling parameter into the STORM optimizer (Cutkosky and Orabona, 2019), effectively resolving the incompatibility of variance-reduction techniques with preconditioned optimization methods, which have demonstrated superior performance in both language modeling and computer vision tasks. It is also worth noting that a similar scaling idea has also been proposed for SVRG in Yin et al. (2024).
In fact, the recent benchmarks of optimizers for LLM pretraining (Wen et al., 2025; Semenov et al., 2025) have empirically demonstrated that matrix-based optimizers (e.g., Shampoo, SOAP and Muon) outperform scalar-based optimizers (e.g., AdamW and Lion), and variance-reduction approaches (e.g., MARS, NadamW (Dozat, 2016)) also yield additional, discernible speedup. Therefore, a natural question arises:
Can we achieve the best of both worlds by combining matrix-based optimizers with variance reduction such as MARS?
It is worth mentioning that in the original MARS work (Yuan et al., 2025), the authors have already proposed MARS-shampoo, which attempts to combine matrix-based optimizer with variance reduction. However, MARS-shampoo displays performances inferior to vector-based variants like MARS-AdamW, leaving open the question of whether variance reduction can be effectively integrated with matrix-based optimizers. In this paper, we give an affirmative answer to this question by introducing MARS-M, a matrix-based MARS optimizer that integrates variance reduction with the Moonlight version (Liu et al., 2025) of Muon. Additionally, we also propose the approximated version of MARS-M as a practical variant, which is also the bridge between variance reduction techniques and traditional preconditioned optimizers.
In summary, our contributions are highlighted as follows:
  • We propose MARS-M, an instantiation of MARS tailored to the Moonlight optimizer. In addition, we develop an approximate version of MARS-M to further accelerate training. We show the connections and differences between MARS-M and MARS-Shampoo. We also show that the approximate version of MARS-M can be seen as a variant of Moonlight with adjusted momentum parameters.
  • We further provide a convergence analysis of MARS-M, and show that it attains a convergence rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, which surpasses the previously established $\mathcal{O}(T^{-1/4})$ rate achieved by Muon (Li and Hong, 2025; Shen et al., 2025; Pethick et al., 2025).
  • Empirically, we evaluate the performances of MARS-M as well as Muon (Moonlight) optimizer for training GPT-2 (Radford et al., 2019) series models on OpenWebText and FineWeb-Edu 100B datasets. The experiment results show that MARS-M can manifest stable improvement on training and validation losses. Moreover, on downstream tasks, MARS-M can also achieve better performances on benchmarks such as Hellaswag (Zellers et al., 2019) and SciQ (Johannes Welbl, 2017) than baseline optimizers. Additionally, MARS-M can also achieve better test accuracy than AdamW and Muon on computer vision tasks.
Notation We use bold capital letters $\mathbf{X},\mathbf{Y},\cdots$ to denote matrices. For a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we use $\|\mathbf{W}\|_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}\mathbf{W}_{ij}}$ to denote its Frobenius norm. In this paper, we use $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ to represent the parameters of the language model at training step $t$, where the training data for each step is a sequence of independent random variables, $\xi_{1},\dots,\xi_{T}\in\mathbf{\Xi}$. For a differentiable objective function $f$, we assume that the expected value of $f(\mathbf{X},\xi_{t})$ given $\mathbf{X}$ is $F(\mathbf{X})$ for all $\mathbf{X}$ and $t$. In this paper, we may explicitly denote the input data, since variance reduction algorithms may utilize both the training data from the current step ($\xi_{t}$) and the previous step ($\xi_{t-1}$) to calculate different gradients for the same parameter. MARS-M is directly inspired by the success of matrix-based optimizers Muon and moonlight.
2 Related Work
Preconditioned gradient methods Preconditioned gradient methods, inspired by Newton’s method, are a powerful class of optimization algorithms widely used in training deep neural networks. To reduce the computational and memory costs of computing and storing the full Hessian matrix, a lot of approximate preconditioners have been proposed, including diagonal approximations (Duchi et al., 2011; Loshchilov and Hutter, 2019; Liu et al., ; Yao et al., 2021) and sketched approximations (Erdogdu and Montanari, 2015; Gonen and Shalev-Shwartz, 2015). Despite these advances, there are only a few practical matrix-based optimization methods for large-scale models. A notable breakthrough was the Shampoo optimizer (Gupta et al., 2018), which demonstrated the practical feasibility of incorporating second-order information in large-scale training. To reduce the number of hyperparameters and computational overhead of Shampoo, Vyas et al. (2024) proposed SOAP, which implements AdamW in the eigenbasis of Shampoo’s preconditioners, significantly improving the training efficiency in terms of both number of iterations and wall clock time. Recently, Muon (Jordan et al., 2024) is proposed by leveraging a Newton–Schulz approximation of the singular value decomposition (SVD) of the gradient momentum, while Moonlight (Liu et al., 2025) extends Muon by introducing weight decay and coupling step sizes across components optimized by AdamW and Muon, thereby improving hyperparameter search efficiency. Building on this line of research, PolarGrad (Lau et al., 2025) introduces a preconditioning framework based on the polar decomposition of the gradient matrix, generalizing the principles of Muon. Scion (Pethick et al., 2025) further unifies Muon and related methods under a linear minimization oracle (LMO) framework, and Gluon (Riabinin et al., 2025) extends both Muon and Scion by incorporating momentum-based generalizations. Nowadays, matrix-based optimizers are becoming increasingly more popular in training industrial large language models, including Kimi K2 (Team et al., 2025) and GLM-4.5 (Zeng et al., 2025).
In order to understand the convergence of Muon, Li and Hong (2025)Shen et al. (2025) and Pethick et al. (2025) proved that Muon has a convergence rate of $\mathcal{O}(T^{-1/4})$. An et al. (2025) and Shen et al. (2025) also analyzed the convergence of Muon without momentum. Kovalev (2025) investigated the convergence rate of Muon with additional assumptions of spectral norm-based Lipschitz smoothness and second-order smoothness, while Li and Hong (2025)Sfyraki and Wang (2025) as well as Sato et al. (2025) analyzed the convergence rate under big batch sizes. Our analysis of MARS-M pushes the theoretical frontier of Muon-based optimization algorithms.
Variance Reduction Methods. The earliest efforts to accelerate stochastic gradient descent (SGD) through variance reduction include SAG (Roux et al., 2012) and SDCA (Shalev-Shwartz and Zhang, 2013). These were soon followed by simpler yet equally effective algorithms such as SVRG (Johnson and Zhang, 2013) and SAGA (Defazio et al., 2014), which achieved the same improved convergence guarantees. Building on this line of work, SARAH (Nguyen et al., 2017a) introduced a biased recursive gradient estimator that reduces memory requirements while retaining optimal complexity bounds for convex optimization. Additionally, variance reduction has also been studied in conjunction with preconditioning in the convex setting (Frangella et al., 2024; Derezinski, 2023). In the non-convex regime, algorithms like SVRG (Allen-Zhu and Yuan, 2016; Reddi et al., 2016) and SARAH (Nguyen et al., 2017b) paved the way for methods such as SPIDER (Fang et al., 2018), which integrates normalized gradient descent (Nesterov, 2013; Hazan et al., 2015) with stochastic path-integral differential estimator, and SNVRG (Zhou et al., 2020), which leverages multiple reference points to enhance variance reduction. SpiderBoost (Wang et al., 2019) further improved SPIDER by permitting large constant step sizes without compromising near-optimal oracle complexity. Subsequently, STORM (Cutkosky and Orabona, 2019) streamlined SPIDER and SNVRG via stochastic recursive momentum, and was later extended into a parameter-free variant, STORM+ (Levy et al., 2021). Recent work has also explored combining variance reduction with adaptive gradient methods. For instance, Adam+ (Liu et al., 2020) reduces variance in Adam by estimating gradients only at extrapolated points, while SuperAdam (Huang et al., 2021) and VRAdam (Li, 2024) integrate variance reduction into AdamW to accelerate convergence. AdaSPIDER (Kavis et al., 2022) extends SPIDER with adaptive step sizes. However, these variance-reduced adaptive optimization algorithms have been evaluated only on relatively simple computer vision and natural language modeling benchmarks with modest model sizes. To our knowledge, MARS (Yuan et al., 2025) is the first method to achieve stellar performance on large language models. This work pointed out that previous algorithms introduced excessive gradient correction and it incorporates scaled gradient correction into adaptive gradient methods, achieving better performances in language modeling and computer vision tasks. MARS-M is built upon MARS, while utilizing matrix-based optimizer muon/moonlight.
3 Preliminaries
In this section, we present the problem setting and relevant preliminaries, including a review of the Muon and MARS optimizers.
In this paper, we consider minimizing an objective function $F(\cdot):\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ as follows:
$\min_{\mathbf{X}}F(\mathbf{X})=\mathbb{E}_{\bm{\xi}\sim\mathcal{D}}[f(\mathbf{X},\bm{\xi})],$ (3.1)
where $f(\mathbf{X},\bm{\xi})$ is possibly nonconvex loss function, $\mathbf{X}\in\mathbb{R}^{m\times n}$ is matrix-variate optimization variable, $\bm{\xi}$ is a random vector (e.g., a training data point) drawn from an unknown data distribution $\mathcal{D}$. We assume the access to the first-order oracle, which returns an unbiased estimator of the gradient $\mathbb{E}[\nabla f(\mathbf{X},\bm{\xi})]=\nabla F(\mathbf{X})$. Throughout the paper, without loss of generality, we assume $m\geq n$.
3.1 Muon
Muon (Jordan et al., 2024) was proposed to exploit the 2D geometric information of model parameter matrices during optimization. Given the momentum $\beta\in(0,1)$ and learning rate $\eta_{t}>0$, the update rule for Muon is as follows:
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (3.2)
$\mathbf{O}_{t}$ $=\text{NewtonSchulz}\left(\mathbf{M}_{t}\right),$ (3.3)
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta_{t}\mathbf{O}_{t}.$ (3.4)
Here Newton-Schulz iteration (Bernstein and Newhouse, 2024) is utilized to approximate $\mathbf{U}_{t}\mathbf{V}_{t}$, where $\mathbf{U}_{t}\bm{\Sigma}_{t}\mathbf{V}_{t}=\mathbf{M}_{t}$ is the Singular Value Decomposition (SVD) of the momentum matrix. In practice, $\widetilde{\mathbf{M}_{t}}=\beta\mathbf{M}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ instead of $\mathbf{M}_{t}$ is usually applied in NewtonSchulz iteration [^1] In Jordan et al. (2024)’s newest implementation, both $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ in $\widetilde{\mathbf{M}_{t}}$ and in (3.2) are substituted with $(1-\beta)\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$. However, in this paper, we just focus on the original algorithm..
It is worth noting that Muon is targeted only for optimizing matrix-like parameters, while AdamW is often applied to optimize vector-like parameters, including embeddings, language model head and RMSNorm. However, a lot of empirical evidences (Yuan et al., 2025; Semenov et al., 2025) found that the original version of Muon does not perform so satisfactorily. Liu et al. (2025) discovered that the inferior performance is due to mismatched updates on the two kinds of parameters (i.e., matrix-like parameters and vector-like parameters). In detail, the RMS norm of AdamW update typically ranges from approximately 0.2 to 0.4 during LLM training. Based on this observation, they proposed a variant of Muon (referred to as “Moonlight” in this paper):
$\mathbf{U}_{t}$ $=\beta\mathbf{U}_{t-1}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (3.5)
$\mathbf{M}_{t}$ $=\beta\mathbf{U}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (3.6)
$\mathbf{O}_{t}$ $=\text{NewtonSchulz}\left(\mathbf{M}_{t}\right),$ (3.7)
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta_{t}(0.2\cdot\mathbf{O}_{t}\cdot\sqrt{\max(m,n)}+\lambda\mathbf{X}_{t}),$ (3.8)
where $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ is the matrix-like parameter to be optimized, and $\lambda>0$ is the decoupled weight decay parameter. Such a corrected version demonstrates great performance in a lot of benchmarks (Semenov et al., 2025; Wen et al., 2025) and successfully applied in training industrial large language models (Team et al., 2025; Zeng et al., 2025).
3.2 MARS
To accelerate the convergence of SGD, a lot of variance reduction techniques have been proposed trying to reduce the variance in update and achieve stabler and faster training.
A typical form for variance reduction is exemplified by STORM (Cutkosky and Orabona, 2019), which introduces gradient correction to the momentum update rule:
$\mathbf{m}_{t}=\beta\mathbf{m}_{t-1}+(1-\beta)\Big[\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{\frac{\beta}{1-\beta}\big(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big)}_{\text{gradient correction}}\Big].$
where $\beta>0$ is momentum parameter. Here the noise brought by the randomness of data can be canceled out by the stochastic gradient difference term $\frac{\beta}{1-\beta}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t}))$. Based on this term, the estimation of the true gradient for the parameter in the last step $\nabla F(\mathbf{x}_{t-1})$ can be transferred to the true gradient for the parameter in the current step $\nabla F(\mathbf{x}_{t})$.
Based on STORM, MARS (Yuan et al., 2025) (See Algorithm 1) introduces a scaling parameter to the gradient correction part such that the corrected gradient is calculated as:
$\bm{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\underbrace{\frac{\beta}{1-\beta}\big(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big)}_{\text{scaled gradient correction}},$ (3.9)
where $\gamma_{t}>0$ is the coefficient for scaling. Moreover, it also introduces gradient clipping on the corrected gradient for stabler training:
$\widetilde{\mathbf{c}}_{t}=\text{Clip}(\mathbf{c}_{t},1)=\begin{cases}\frac{\mathbf{c}_{t}}{\|\mathbf{c}_{t}\|_{2}}&\text{if }\|\mathbf{c}_{t}\|_{2}>1,\\ \mathbf{c}_{t}&\text{otherwise}.\end{cases}$ (3.10)
For the ease of exposure, the MARS algorithm is summarized in Algorithm 1, where line 7 is the general formula for adaptive gradient methods such as AdamW (Loshchilov and Hutter, 2019), Lion (Chen et al., 2023) and Shampoo (Gupta et al., 2018), with approximated Hessian matrix $\mathbf{H}_{t}$. As will be seen later, our work can be seen as a tailored version of MARS for Muon/moonlight.
Algorithm 1 MARS
1:input: $\mathbf{x}_{0},\beta,\{\gamma_{t}\},\{\eta_{t}\}$
2: Set $\mathbf{m}_{0}\leftarrow\mathbf{0}$ and $\mathbf{x}_{1}\leftarrow\mathbf{x}_{0}$
3:for $t=1,$ to $T$ do
4:  Sample $\bm{\xi}_{t}$ and let $\mathbf{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\gamma_{t}\frac{\beta}{1-\beta}\big(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big)$
5:  if $\|\mathbf{c}_{t}\|_{2}>1$, then $\tilde{\mathbf{c}}_{t}=\frac{\mathbf{c}_{t}}{\|\mathbf{c}_{t}\|_{2}}$ else $\widetilde{\mathbf{c}}_{t}=\mathbf{c}_{t}$
6:  $\mathbf{m}_{t}=\beta\mathbf{m}_{t-1}+(1-\beta)\tilde{\mathbf{c}}_{t}$
7:  $\mathbf{x}_{t+1}=\arg\min_{\mathbf{x}}\left\{\eta_{t}\left\langle\mathbf{m}_{t},\mathbf{x}\right\rangle+\frac{1}{2}\|\mathbf{x}-\mathbf{x}_{t}\|_{\mathbf{H}_{t}}^{2}\right\}$
8:end for
4 Method
In this section, we propose the MARS-M optimizer and establish its convergence guarantees.
4.1 MARS-M
We present MARS-M, which is an instantiation of MARS by substituting the update rule (3.4) with (3.8), and the resulted algorithm is shown in Algorithm 2. MARS-M is an instantiation of MARS tailored to the Moonlight optimizer.
It is worth noting that MARS-M is similar to MARS-Shampoo in (Yuan et al., 2025), which is another instantiation of MARS to a simplified version of Shampoo. In detail, the update rule in MARS-Shampoo is instantiated as:
$\mathbf{U}_{t},\mathbf{\Sigma}_{t},\mathbf{V}_{t}$ $=\text{SVD}(\mathbf{M}_{t}),$
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t}\mathbf{V}_{t}^{\top}.$ (4.1)
In practice, SVD computation can be substituted by approximate numerical methods such as Newton iteration (Lakić, 1998; Higham, 2008; Anil et al., 2020; Bernstein and Newhouse, 2024) and Newton-Schulz iteration (Schulz, 1933; Higham, 2008). The main differences between MARS-M and MARS-Shampoo are that MARS-M scales $\mathbf{O}_{t}\approx\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ by $0.2\cdot\sqrt{\max(m,n)}$ and incorporates weight decay $\lambda\mathbf{X}_{t}$, both of which are essential for achieving superior performance according to Liu et al. (2025).
Algorithm 2 MARS-M
1:input: $\mathbf{X}_{0}\in\mathbb{R}^{m\times n},\lambda,\beta,\{\gamma_{t}\},\{\eta_{t}\}$
2: Set $\mathbf{M}_{0}\leftarrow\mathbf{0}$ and $\mathbf{X}_{1}\leftarrow\mathbf{X}_{0}$
3:for $t=1,$ to $T$ do
4:  sample $\bm{\xi}_{t}$ and let $\mathbf{C}_{t}=\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})+\gamma_{t}(\frac{\beta}{1-\beta})\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t})\big)$
5:  $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\text{Clip}(\mathbf{C}_{t},1)$
6:  $\mathbf{O}_{t}=\text{NewtonSchulz}(\mathbf{M}_{t})$
7:  $\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}(0.2\cdot\mathbf{O}_{t}\cdot\sqrt{\max(m,n)}+\lambda\mathbf{X}_{t})$
8:end for
4.2 Approximated MARS-M
Since calculating the stochastic gradients twice requires much more computation, in practice, approximate version of MARS is usually utilized where $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t})$ is replaced by $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})$. If clipping on $\mathbf{C}_{t}$ is ignored, lines 4 to 5 in Algorithm 2 can be approximated by:
$\mathbf{C}_{t}$ $=\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})+\gamma_{t}\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.2)
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{C}_{t}.$
With some calculation, it can be shown that the approximate version of MARS-M without clipping is equivalent to the following Muon/Moonlight-style update rule:
$\mathbf{U}_{t}$ $=\beta\mathbf{U}_{t-1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{(1-\gamma_{t})(1-\beta)}{\beta}}\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (4.3)
$\mathbf{M}_{t}$ $=\beta\mathbf{U}_{t}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (4.4)
$\mathbf{O}_{t}$ $=\text{NewtonSchulz}\left(\mathbf{M}_{t}\right),$
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta_{t}(0.2\cdot\mathbf{O}_{t}\cdot\sqrt{\max(m,n)}+\lambda\mathbf{X}_{t}).$
By comparing (4.3), (4.4) with (3.5), (3.6), we can see that approximate MARS-M can be seen as a variant of Moonlight with adjusted momentum parameters $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(1-\gamma_{t})(1-\beta)/\beta$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}$ (in contrast to 1’s in moonlight).
From another perspective, we note that (3.5) and (3.6) in the original Moonlight algorithm can be equivalently expressed as:
$\mathbf{C}_{t}$ $={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{1}{1-\beta}}\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}\cdot\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.5)
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{C}_{t}.$~ (4.6)
Comparing (4.2) with (4.5), it can be seen that approximate version of MARS-M is Moonlight with stochastic gradient $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ being scaled down by a factor of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\beta)$ and setting $\gamma_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}$. The experiments in Section 5 empirically demonstrate that this difference in momentum parameters leads to a noticeable impact on the loss.
4.3 Convergence Analysis
To analyze the convergence of MARS-M, similar to Yuan et al. (2025), we first make the following assumptions:
Assumption 4.1 (Bounded Variance).
We assume that the variance of gradient estimator is bounded by $\sigma^{2}$. i.e., for any noise $\bm{\xi}$, parameter $\mathbf{X}$, and $\nabla F(\mathbf{X})=\mathbb{E}[\nabla f(\mathbf{X},\bm{\xi})]$, there exists a positive $\sigma$ such that:
$\mathbb{E}\big[\|\nabla f(\mathbf{X},\bm{\xi})-\nabla F(\mathbf{X})\|_{F}^{2}\big]\leq\sigma^{2}.$ (4.7)
Assumption 4.2 ($L$-Smoothness).
We assume that for arbitrary $\bm{\xi}$, $f(\mathbf{X},\bm{\xi})$ is $L$-smooth:
$\|\nabla f(\mathbf{X},\bm{\xi})-\nabla f(\mathbf{Y},\bm{\xi})\|_{F}\leq L\cdot\|\mathbf{X}-\mathbf{Y}\|_{F},~\forall\mathbf{X},\mathbf{Y}.$ (4.8)
Both assumptions are standard in the literature (Cutkosky and Orabona, 2019; Yuan et al., 2025). We have the following theorem, which guarantees the convergence of MARS-M in Algorithm 2.
Theorem 4.3.
In Algorithm 2, under Assumptions 4.1 and 4.2, when choosing $\lambda=0$, $\eta_{t}=(s+t)^{-2/3},s\geq 2$, suppose $\beta_{t+1}=1-2\eta_{t}$, then for $\forall T\geq s$, it holds that
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})\|_{F}\leq\frac{2\sqrt{2LG}}{T^{1/3}}+\frac{2\sqrt{2LB\log(s+T)}}{T^{1/3}}+\frac{2G}{T^{1/3}}+\frac{2B}{T^{1/3}}\log(s+T)-\frac{\sqrt{2}}{4LT^{1/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
where $G=F(\mathbf{X}_{1})-\min_{\mathbf{X}}F(\mathbf{X})+\frac{\sqrt{2}s^{1/3}\sigma^{2}}{4L}+\frac{3Ln}{2s^{1/3}}$, $B=\Big(\frac{2\sqrt{2}\sigma^{2}}{L}+\frac{3\sqrt{2}Ln}{2}\Big)$ and $\widetilde{M}_{t+1}$ is a non-negative value defined in (C.1).
We defer the proof of the above theorem to Appendix C. It is worth noting that time-varying momentum parameter $\beta_{t}$ is required for theoretical analysis, while it can be chosen as a constant in practice. Theorem 4.3 suggests that MARS-M can achieve a non-asymptotic convergence rate of $\mathcal{O}(T^{-1/3}\log(T))$, the same rate as general MARS as proved in Yuan et al. (2025). As a comparison, Li and Hong (2025)Shen et al. (2025) and Pethick et al. (2025) proved that Muon can achieve only $\mathcal{O}(T^{-1/4})$ convergence rate.
To sum up, Muon-M constitutes a new addition to the MARS family, broadening the framework to encompass more matrix-based optimizers with strong theoretical guarantees.
5 Experiments
5.1 LLM Experiments
We evaluate the performance of our algorithm with baseline algorithms [^2] For training efficiency, we use approximated MARS-M for LLM experiments, since the performances of exact version and approximate version are similar as discussed in Yuan et al. (2025). And we compare their differences in computer vision experiments., including Moonlight and AdamW, on language modeling tasks based on the nanoGPT (Karpathy, 2022) architecture and GPT-2 (Radford et al., 2019) model. We implement experiments with OpenWebText (Gokaslan et al., 2019) and FineWeb-Edu 100B (Lozhkov et al., 2024) dataset. For OpenWebText, the training and validation datasets contain around $9$ billion and $4.4$ million tokens, respectively; while the training and validation sets of FineWeb-Edu 100B are with 100 billion and 0.1 billion tokens. The experiments are conducted on scales of small (125M parameters), medium (355M parameters), large (770M parameters) as well as XL (1.5B parameters). Moreover, we disabled biases, and set the Dropout rate (Srivastava et al., 2014) to $0.0$. For experiments with small models, we utilize 16 NVIDIA H800 GPUs; and for models with larger sizes, we implement with 32 NVIDIA H800. Other hyper-parameters of training are listed in Appendix A.
5.2 Experiment Results
We show the zoomed-in curves for training and validation losses for large-size models trained on OpenWebText and XL-size models trained on FineWeb-Edu 100B datasets in Figures 3 and 1. Moreover, we also show the curves for XL-size models for all the training process in Figure 2. Other curves can be found in Appendix B. We also add the curves for the experiments trained with AdamW in Yuan et al. (2025), where the maximum learning rate is 6e46e-4 for XL-size model and 2e42e-4 for large-size model. It can be observed that the experiments trained with MARS-M display steady improvement on both training and validation losses over Moonlight. Additionally, although the loss for AdamW is lower in the middle of training due to a smaller maximum learning rate, model trained with MARS-M achieves lower losses than that of AdamW in the final phase. Since the learning rates for all the experiments trained with baseline optimizers results from grid search, it can be concluded that MARS-M can indeed improve the performance of large language model training.
Figure 1: The training and validation loss of XL-size models (1.5B) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 2: The training and validation loss of XL-size models (1.5B) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 3: The training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
Furthermore, We also evaluate 0-shot and 2-shot performances of these optimizers on benchmarks including ARC (Yadav et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SciQ (Johannes Welbl, 2017), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), based on lm-evaluation-harness codebase (Gao et al., 2024). We only display the 2-shot performances for XL models in Table 1, and postpone other results in the Appendix B. It can be observed that MARS-M outperforms Moonlight on most of the benchmarks, showing that our algorithm can enhance the performance of pre-trained large language models on a wide range of downstream tasks.
Table 1: The evaluation results of XL models pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
Muon (Moonlight) 40.87 71.21 55.80 24.89 42.40 74.05 91.90 57.54 57.33
MARS-M ($\gamma=0.01$) 41.64 73.57 57.01 24.75 39.60 73.99 92.30 58.96 57.73
MARS-M ($\gamma=0.025$) 42.66 72.10 57.09 25.57 39.60 74.21 92.30 58.09 57.70
5.3 Computer Vision Experiments
Besides language model task, we also conduct experiments on computer vision tasks with these optimizers on CIFAR-10 dataset (Krizhevsky et al., 2009) and ResNet-18 model (He et al., 2016). We set $\gamma=0.025$ for MARS-M and do grid search over $\{10^{-4},\cdots,10^{-2}\}$ for the learning rate. We set $\beta=0.95$ for all the optimizers. All the experiments are done with training batch size 128 on 1 NVIDIA A6000 GPU, with a total of 200 training epochs. Additionally, we use MultiStepLR scheduler so that the learning rate will decrease to 10% at 100th step and 1% at 150th step. The test loss and accuracy for MARS-M and MARS-M approximate, as well as Moonlight optimizer, are shown in Figure 4. It can be seen that MARS-M achieve better test accuracy and lower test loss over Moonlight and MARS-M-approx, validating the effect of variance reduction.
Figure 4: The test loss and test accuracy for different optimizers on CIFAR-10 dataset.
5.4 Ablation Study
[…]
We also implement ablation study on $\gamma$ in MARS-M to unearth the effect of scale of gradient correction on the performance. The experiments are done on small models trained with OpenWebText datasets, and the results are shown in Figure 5. It can be observed that the scale on gradient correction should be much smaller than 1, as usually taken in most of the variance reduction approaches (Cutkosky and Orabona, 2019; Nguyen et al., 2017a; Fang et al., 2018). However, the training and validation losses for experiments with $\gamma\in[0.005,0.025]$ are very close, showing that a small enough $\gamma$ is relatively robust to variation.
6 Conclusion
In this work, we have introduced MARS-M, a new algorithm that integrates MARS-based variance reduction with Muon/Moonlight to accelerate large language model training. Our algorithm integrates the advantages of variance reduction and matrix-based optimization. Through empirical experiments on language model and computer vision tasks, we demonstrate that models trained with MARS-M gain steady edges on loss and evaluation on downstream tasks over Moonlight baseline. In summary, our algorithm demonstrates an effective pathway for improving state-of-the-art matrix-based optimizers using variance reduction techniques.
Appendix A Hyper-parameter Settings
For training parameters, for experiments with either OpenWebText or FineWeb-Edu 100B datasets, we do a grid search over learning rates between $\{5e-4,1e-3,3e-3,5e-3,6e-3,1e-2\}$ for Moonlight optimizer. We just apply the same learning rate for MARS-M optimizer. Since Moonlight and MARS-M optimizers are only designed for matrix parameters, we optimize the vector-like parameters and embeddings with AdamW with the same learning rate. We list the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL) parameters in Table 2. Moreover, the general hyperparameters used across all experiments are summarized in Table 3, while Table 4 display the training hyperparameters for models with different sizes.
Table 2: Architecture hyperparameters for GPT-2 series models (Radford et al., 2019).
Model #Param #Layer $n_{\text{head}}$ $d_{\text{emb}}$
GPT-2 small 125M 12 12 768
GPT-2 medium 355M 24 16 1024
GPT-2 large 770M 36 20 1280
GPT-2 XL 1.5B 48 25 1600
Table 3: General hyper-parameters for the experiments.
Hyper-parameter Value
Steps 100,000
Batch size in total 480
Context length 1024
Gradient clipping threshold 1.0
Dropout 0.0
Learning rate schedule Cosine
Warm-up steps 2000
Base seed 5000
Table 4: Hyper-parameters for GPT-2 experiments for different datasets. We use $\beta=0.95$ for the optimizers.
Hyper-parameter GPT-2 Size OpenWebText FineWebEdu 100B
Max learning rate small (125M) 6e-3 1e-2
medium (355M) 5e-3 5e-3
large (770M) 5e-3 5e-3
XL (1.5B) - 3e-3
Min learning rate small (125M) 3e-5 3e-5
medium (355M) 6e-5 6e-5
large (770M) 1e-5 1e-5
XL (1.5B) - 1e-5
Appendix B Additional Experiments
We just display the curves for training and validation losses for small, medium and large models on the FineWeb-Edu 100B and OpenWebText datasets in Figures 6-8 and Figures 9-11, respectively. And the 0-shot and 2-shot evaluation results for these models in Tables 5-11 and Tables 12-17, respectively. The results of AdamW are taken from Yuan et al. (2025). According to these figures and tables, it can be observed that MARS-M can achieve better performances than AdamW and Moonlight optimizers in most of these cases.
Figure 6: The training and validation loss of small-size models (125M) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 7: The training and validation loss of medium-size models (355M) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 8: The training and validation loss of large-size models (770M) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 9: The training and validation loss of small-size models (125M) trained with different optimizers on the OpenWebText dataset.
Figure 10: The training and validation loss of medium-size models (355M) trained with different optimizers on the OpenWebText dataset.
Figure 11: The training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
Table 5: The evaluation results of small models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 26.54 51.43 36.26 24.49 30.60 64.53 71.50 50.36 44.46
Muon (Moonlight) 26.88 52.48 37.58 23.12 32.80 64.85 71.20 50.67 44.95
MARS-M ($\gamma=0.01$) 26.28 51.47 37.77 23.26 31.80 65.94 70.90 52.01 44.93
MARS-M ($\gamma=0.025$) 27.82 52.57 37.74 26.01 32.00 65.23 71.90 51.14 45.55
Table 6: The evaluation results of small models pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
Muon (Moonlight) 28.41 57.62 37.00 25.83 30.00 64.20 82.90 51.38 47.17
MARS-M ($\gamma=0.01$) 27.82 58.16 37.50 25.66 31.60 65.02 86.40 53.51 48.21
MARS-M ($\gamma=0.025$) 28.41 58.42 37.06 26.46 31.20 65.34 82.10 50.75 47.47
Table 7: The evaluation results of medium models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
MARS 32.68 59.34 45.05 24.98 35.40 68.88 76.10 54.85 49.66
Muon (Moonlight) 32.17 58.92 45.46 23.28 34.80 69.64 76.80 54.14 49.40
MARS-M ($\gamma=0.01$) 30.55 59.60 46.11 25.13 31.80 69.91 77.50 54.06 49.33
MARS-M ($\gamma=0.025$) 32.25 60.65 45.92 24.52 33.80 67.90 78.30 56.20 49.94
Table 8: The evaluation results of medium models pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 33.53 66.46 45.02 25.32 35.40 69.10 86.70 55.80 52.17
Muon (Moonlight) 33.45 66.33 45.48 24.73 35.40 69.21 89.00 54.06 52.21
MARS-M ($\gamma=0.01$) 34.73 66.12 46.23 24.56 34.00 70.40 88.70 53.83 52.32
MARS-M ($\gamma=0.025$) 35.24 67.30 45.82 24.98 33.40 68.28 89.10 54.93 52.38
Table 9: The evaluation results of large models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 34.64 63.22 50.10 25.45 38.60 70.89 80.60 57.54 52.63
Muon (Moonlight) 34.30 63.34 51.93 24.18 38.80 72.31 81.50 57.77 53.07
MARS-M ($\gamma=0.01$) 36.18 64.52 52.29 25.89 37.60 71.65 82.30 56.59 53.45
MARS-M ($\gamma=0.025$) 31.74 62.84 51.91 24.35 35.80 70.35 79.70 54.85 51.45
Table 10: The evaluation results of large models pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 38.05 70.29 50.30 26.87 38.20 70.46 92.10 55.80 55.26
Muon (Moonlight) 37.03 70.08 51.57 25.32 38.20 72.03 91.00 56.43 55.21
MARS-M ($\gamma=0.01$) 39.08 70.16 52.07 24.90 37.60 72.25 93.30 56.91 55.78
MARS-M ($\gamma=0.025$) 37.03 72.01 52.22 25.10 39.40 71.22 90.90 55.64 55.41
Table 11: The evaluation results of XL models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 38.40 68.22 53.93 25.47 39.00 72.69 85.30 54.78 54.72
Muon (Moonlight) 39.08 67.30 56.37 25.32 43.20 73.78 83.50 58.41 55.87
MARS-M ($\gamma=0.01$) 38.74 68.14 56.83 24.68 39.80 73.23 84.70 57.22 55.42
MARS-M ($\gamma=0.025$) 38.31 68.56 57.12 23.57 41.40 73.72 85.30 56.75 55.59
Table 12: The evaluation results of small models pre-trained using the OpenWebText dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 22.27 41.37 31.73 22.97 27.80 63.00 67.50 52.01 41.08
Muon (Moonlight) 23.81 41.67 33.19 22.83 28.60 63.06 66.50 51.38 41.38
MARS-M ($\gamma=0.01$) 23.72 40.66 33.36 22.97 27.80 74.80 65.10 51.38 42.47
MARS-M ($\gamma=0.025$) 23.04 40.49 33.44 22.93 30.20 63.22 66.40 50.83 41.32
Table 13: The evaluation results of small models pre-trained using the OpenWebText dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
Muon (Moonlight) 24.23 43.18 33.20 25.00 26.80 62.79 75.20 53.91 43.04
MARS-M ($\gamma=0.01$) 24.49 43.52 33.47 25.35 26.20 62.02 73.50 51.46 42.50
MARS-M ($\gamma=0.025$) 23.38 43.43 33.39 23.99 27.40 62.40 77.40 49.41 42.60
Table 14: The evaluation results of medium models pre-trained using the OpenWebText dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 23.98 43.43 37.76 22.80 27.20 65.56 67.60 52.49 42.60
Muon (Moonlight) 25.60 44.99 40.00 23.20 31.20 65.94 72.00 52.25 44.40
MARS-M ($\gamma=0.01$) 24.49 44.49 40.22 23.34 33.40 67.08 68.01 54.46 44.44
MARS-M ($\gamma=0.025$) 24.57 46.42 40.21 23.28 32.00 66.27 70.00 52.64 44.42
Table 15: The evaluation results of medium models pre-trained using the OpenWebText dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
Muon (Moonlight) 27.13 48.82 39.77 24.93 31.40 67.03 81.30 52.80 46.65
MARS-M ($\gamma=0.01$) 26.79 50.17 40.87 25.42 28.80 67.14 81.60 52.64 46.68
MARS-M ($\gamma=0.025$) 26.37 48.82 40.53 24.68 31.60 65.78 82.40 53.67 46.73
Table 16: The evaluation results of large models pre-trained using the OpenWebText dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 26.19 46.30 41.70 23.10 31.40 68.12 72.80 51.46 45.13
Muon (Moonlight) 27.30 47.85 45.56 23.85 30.80 68.93 72.50 55.64 46.55
MARS-M ($\gamma=0.01$) 26.88 49.37 45.74 24.41 32.00 69.37 74.30 53.43 46.94
MARS-M ($\gamma=0.025$) 27.05 48.48 45.19 23.59 33.20 68.82 72.50 57.30 47.02
Table 17: The evaluation results of large models pre-trained using the OpenWebText dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: WG = WinoGrande.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
Muon (Moonlight) 27.90 53.45 45.27 25.77 32.80 69.53 84.20 56.12 49.38
MARS-M ($\gamma=0.01$) 27.99 53.96 45.94 24.76 31.40 68.82 86.40 56.27 49.44
MARS-M ($\gamma=0.025$) 27.47 53.07 45.33 25.10 31.80 69.21 85.50 55.49 49.12
Appendix C Proof of Theorem 4.3
We have the following lemma from Yuan et al. (2025):
Lemma C.1 (Yuan et al. 2025).
In Algorithm 2. Under Assumption 4.1 and 4.2, if $1\geq\beta_{t+1}\geq 0$, $\forall t$, denote $\bm{\Delta}_{t}:=\nabla f(\mathbf{X}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{X}_{t},\bm{\xi}_{t+1})$, under approximate choice of $\gamma_{t+1}$:
$\gamma_{t+1}=1-\frac{\Big(G_{t+1}+\beta_{t+1}\big(\mathbb{E}\|\bm{\Delta}_{t}\|_{F}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{F}^{2}\big)\Big)}{\beta_{t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{F}^{2}}=\frac{\beta_{t+1}\|\mathbb{E}\bm{\Delta}_{t}\|_{F}^{2}-G_{t+1}}{\beta_{t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{F}^{2}},$
we have
$\mathbb{E}\|\nabla F(\mathbf{X}_{t+1})-\mathbf{M}_{t+1}\|_{F}^{2}$ $\leq\beta_{t+1}^{2}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}-\widetilde{M}_{t+1},$
where
$\widetilde{M}_{t+1}:=\mathbb{E}\|\bm{\Delta}_{t}\|_{F}^{2}\Bigg(A_{t+1}^{2}-\Big(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1}\Big)^{2}\Bigg),$ (C.1)
$A_{t+1}:=\frac{G_{t+1}+\beta_{t+1}\text{tr}\big(\operatorname{{\rm Var}}\big(\bm{\Delta}_{t}\big)\big)}{\mathbb{E}\|\bm{\Delta}_{t}\|_{F}^{2}},$
and
$G_{t+1}:=(1-\beta_{t+1})\mathbb{E}\bigg\langle\bm{\Delta}_{t},\nabla f(\mathbf{X}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{X}_{t+1})\bigg\rangle+\beta_{t+1}\mathbb{E}\bigg\langle\nabla F(\mathbf{X}_{t+})-\nabla F(\mathbf{X}_{t}),F(\mathbf{X}_{t})-\mathbf{M}_{t}\bigg\rangle.$
We also need the following lemmas, which can be seen as an extension of Lemma C.3 in Yuan et al. (2025).
Lemma C.2.
For Algorithm 2, under Assumption 4.2, we have the following inequality hold:
$F(\mathbf{X}_{t+1})\leq F(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}}{\rho}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}\rho}{4}\|\mathbf{O}_{t}\|_{F}^{2}+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2},$
where $\rho>0$ is a an arbitrarily chosen parameter.
Proof.
According to Assumption 4.2, we have the upper bound for the function value:
$F(\mathbf{X}_{t+1})$ $\leq F(\mathbf{X}_{t})+\langle\nabla F(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$=F(\mathbf{X}_{t})+\langle\nabla F(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$=F(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
Here $\mathbf{O}_{t}=\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ is defined in (4.1). According to the definition of $\mathbf{O}_{t}$, we have:
$-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F},$
where the last inequality follows from $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$. Moreover, by AM-GM inequality, it holds that
$\eta_{t}\langle\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle\leq\eta_{t}\left(\frac{1}{\rho}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho}{4}\|\mathbf{O}_{t}\|_{F}^{2}\right)$
for some constant $\rho>0$ . Putting all pieces together, we can obtain
$F(\mathbf{X}_{t+1})\leq F(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}}{\rho}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}\rho}{4}\|\mathbf{O}_{t}\|_{F}^{2}+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
This completes the proof. ∎
Lemma C.3.
Let $\eta_{t}=(s+t)^{-2/3},s\geq 1$, $\forall t\geq 1$. Then $\eta_{t}^{-1}-\eta_{t-1}^{-1}\leq\eta_{t}^{1/2}$, $\forall t\geq 1$.
Proof.
By the definition of $\eta_{t}$, it holds that
$\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}=(s+t)^{2/3}-(s+t-1)^{2/3}\leq\frac{2}{3(s+t-1)^{1/3}}\leq\eta_{t}^{1/2},$
where the first inequality follows by the concavity of $h(\mathbf{x})=x^{2/3}$. This finishes the proof. ∎
Now we’re ready to prove Theorem 4.3.
Proof of Theorem 4.3.
 [^3] Here we just ignore the factor $0.2\cdot\sqrt{\max(m,n)}$ since such a factor can be integrated into $\eta_{t}$.
First, we define the Lyapunov function as
$\Phi_{t}=\mathbb{E}\Big[F(\mathbf{X}_{t})+\frac{\rho_{t}}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}\Big],\quad\forall t\geq 1.$
where $\rho_{t}=4\sqrt{2}L\sqrt{\eta_{t}}$. Then we calculate the difference between two consecutive Lyapunov functions as:
$\Phi_{t+1}-\Phi_{t}$ $=\underbrace{\mathbb{E}[F(\mathbf{X}_{t+1})-F(\mathbf{X}_{t})]}_{\mbox{$I_{1}$}}$
$\qquad+\underbrace{\mathbb{E}\Bigg[\frac{\rho_{t+1}}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{X}_{t+1})-\mathbf{M}_{t+1}\|_{F}^{2}-\frac{\rho_{t}}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}\Bigg]}_{\mbox{$I_{2}$}}.$ (C.2)
For $I_{1}$, we use Lemma C.2 to obtain
$I_{1}\leq\mathbb{E}\Big[-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}}{\rho_{t}}\cdot\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}\rho_{t}}{4}\|\mathbf{O}_{t}\|_{F}^{2}+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\Big].$ (C.3)
For $I_{2}$, we use Lemma C.1 to obtain
$I_{2}$ $=\mathbb{E}\Big[\frac{\rho_{t+1}}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{X}_{t+1})-\mathbf{M}_{t+1}\|_{F}^{2}-\frac{\rho_{t}}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}\Big]$
$\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}\beta_{t+1}^{2}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}(1-\beta_{t+1})^{2}\sigma^{2}}{8L^{2}\eta_{t}}$
$\qquad-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1}$
$\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1},$ (C.4)
where the first inequality is due to $\rho_{t+1}\leq\rho_{t}$, and the last inequality follows from the definition that $\beta_{t+1}=1-2\eta_{t}$. Further, for the first term on the right hand side, we have
$\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)$ $\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)=\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{1-2\eta_{t}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)$
$=\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}-2\Big).$
From Lemma C.3, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\leq\eta_{t}^{1/2}$. Since $\rho_{t}^{2}=32L^{2}\eta_{t}$, we obtain [^4] As long as $\eta_{t}\leq 1/2$, $1\geq\beta_{t+1}=1-2\eta_{t}>=0$ is satisfied.
$\frac{\rho_{t}}{16L^{2}}\cdot\Big(\frac{\beta_{t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(\eta_{t}^{1/2}-c\Big)\leq\frac{\rho_{t}}{16L^{2}}\cdot\Big(1-c\Big)=-2\eta_{t}\rho_{t}^{-1}.$ (C.5)
Bringing (C.5) into (C.4), we arrive at the upper bound for $I_{2}$:
$I_{2}\leq-\frac{2\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1}.$ (C.6)
Now combining (C), (C.3) and (C.6), we derive
$\Phi_{t+1}-\Phi_{t}$ $\leq-\frac{\eta_{t}}{\rho_{t}}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}-\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}+\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\frac{\rho_{t}c^{2}\eta_{t}\sigma^{2}}{8L^{2}}$
$\qquad-\frac{\rho_{t}}{16L^{2}\eta_{t}}\widetilde{M}_{t+1}+\frac{\eta_{t}\rho_{t}}{4}\|\mathbf{O}_{t}\|_{F}^{2}+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}.$ (C.7)
Moreover, according to the definition of $\mathbf{O}_{t}$, we have
$\|\mathbf{O}_{t}\|_{F}^{2}\leq\sum_{i=1}^{n}1=n.$
Therefore,
$\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$ $=\frac{\rho_{t}}{8\eta_{t}}\cdot\mathbb{E}\|\eta_{t}\mathbf{O}_{t}\|_{F}^{2}=\frac{\rho_{t}\eta_{t}}{8}\cdot\mathbb{E}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{\rho_{t}\eta_{t}n}{8},$
$\frac{\eta_{t}\rho_{t}}{4}\|\mathbf{O}_{t}\|_{F}^{2}$ $\leq\frac{\rho_{t}\eta_{t}n}{4},$
$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$ $\leq\frac{L\eta_{t}^{2}n}{2}.$
Taking a telescoping sum for $t=1,\cdots,T$ in (C.7), and applying $\rho_{t}=4\sqrt{2}L\sqrt{\eta_{t}}$ give
$\sum_{t=1}^{T}\Big(\frac{\sqrt{\eta_{t}}}{4\sqrt{2}L}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}\Big)$
$\leq\Phi_{1}-\Phi_{T+1}+\frac{\sigma^{2}c^{2}}{8L^{2}}\sum_{t=1}^{T}\rho_{t}\eta_{t}+\sum_{t=1}^{T}\big(\frac{3\rho_{t}\eta_{t}n}{8}+\frac{L\eta_{t}^{2}n}{2}\big)-\sum_{t=1}^{T}\frac{\sqrt{2}}{4L\sqrt{\eta_{t}}}\widetilde{M}_{t+1}$
$\leq\Phi_{1}-\Phi_{T+1}+\Big(\frac{2\sqrt{2}\sigma^{2}}{L}+\frac{3\sqrt{2}Ln}{2}\Big)\sum_{t=1}^{T}\frac{1}{s+t}+\frac{Ln}{2}\sum_{t=1}^{T}\frac{1}{(s+t)^{4/3}}-\sum_{t=1}^{T}\frac{\sqrt{2}}{4L\sqrt{\eta_{t}}}\widetilde{M}_{t+1}$
$\leq\Phi_{1}-\Phi_{T+1}+B\cdot\log(s+T)+\frac{3Ln}{2s^{1/3}}-\sum_{t=1}^{T}\frac{\sqrt{2}}{4L\sqrt{\eta_{t}}}\widetilde{M}_{t+1},$
where $B=\Big(\frac{2\sqrt{2}\sigma^{2}}{L}+\frac{3\sqrt{2}Ln}{2}\Big)$. By the definition of $\Phi_{t}$, we have $\Phi_{T+1}\geq F(\mathbf{X}_{T+1})\geq\min_{\mathbf{X}}F(\mathbf{X})$. And for $\Phi_{1}$,
$\Phi_{1}$ $=\mathbb{E}\Big[F(\mathbf{X}_{1})+\frac{\rho_{1}s^{2/3}}{16L^{2}}\cdot\|\nabla F(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}\Big]$
$\leq F(\mathbf{X}_{1})+\frac{\sqrt{2}s^{1/3}}{4L}\cdot\mathbb{E}[\|\nabla F(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1},\bm{\xi}_{1})\|_{F}^{2}]$
$\leq F(\mathbf{X}_{1})+\frac{\sqrt{2}s^{1/3}\sigma^{2}}{4L}.$
Consequently, defining $G=F(\mathbf{X}_{1})-\min_{\mathbf{X}}F(\mathbf{X})+\frac{\sqrt{2}s^{1/3}\sigma^{2}}{4L}+\frac{3Ln}{2s^{1/3}}$, the following inequality holds:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}$ $\leq\frac{4\sqrt{2}LG}{T\sqrt{\eta_{T}}}+\frac{4\sqrt{2}LB}{T\sqrt{\eta_{T}}}\log(s+T)-\frac{2}{T\sqrt{\eta_{T}}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}$
$\leq\frac{8LG}{T^{2/3}}+\frac{8LB}{T^{2/3}}\log(s+T)-\frac{2}{T^{2/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}},$
where the last inequality holds when $T\geq s$. And it also holds that
$\frac{1}{T}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$ $\leq\frac{G}{T}+\frac{B}{T}\log(s+T)-\frac{\sqrt{2}}{4LT}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
Therefore,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$ $\leq\frac{G}{T\eta_{T}}+\frac{B}{T\eta_{T}}\log(s+T)-\frac{\sqrt{2}}{4LT\eta_{T}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}$
$\leq\frac{G(s+T)^{2/3}}{T}+\frac{B(s+T)^{2/3}}{T}\log(s+T)-\frac{\sqrt{2}(s+T)^{2/3}}{4LT}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}$
$\leq\frac{2G}{T^{1/3}}+\frac{2B}{T^{1/3}}\log(s+T)-\frac{\sqrt{2}}{4LT^{1/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
Finally, by triangle inequality, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})\|_{F}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\frac{1}{T}\sum_{t=1}^{T}\sqrt{\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\sqrt{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\frac{2\sqrt{2LG}}{T^{1/3}}+\frac{2\sqrt{2LB\log(s+T)}}{T^{1/3}}+\frac{2G}{T^{1/3}}+\frac{2B}{T^{1/3}}\log(s+T)-\frac{\sqrt{2}}{4LT^{1/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
This completes the proof. ∎